A php 300 per share dividend was paid by a certain stock. The dividend is estimated to grow at 4.75% per year until further notice. What is the price of the stock today if the required return is 8.75%


1
Expert's answer
2021-12-13T08:35:46-0500

The price of a stock is calculated as

"P=\dfrac{D_1}{r-g}"

Therefore, if a stock paid 300 in dividends and the dividends are expected to grow at the rate of 4.75% per year, the price of the stock if the rate of return is 8.75% is equal to

"P=\dfrac{300}{0.0875-0.0475}\\[0.3cm]
P=\boxed{7500}"